OpenStudy (anonymous):

Solve for r: C = 2(pi)r.

OpenStudy (anonymous):

divide both sides by \[2 \pi\]

OpenStudy (anonymous):

theres no number on both sides how would that work ?

OpenStudy (anonymous):

to solve for r, you have to isolate it by itself. it does't matter whether there is a number or not.

OpenStudy (anonymous):

ok so 2(3.14)r ? and i do it again for c ?

OpenStudy (anonymous):

\[C=2\pi r\]; to solve for r, isolate it by dividing both sides with... \[C \div 2\pi=2\pi r \div 2\pi\]

OpenStudy (anonymous):

so would it be r=c/(pi)

OpenStudy (anonymous):

don't forget the 2